# Generalized Linear Bandits with Limited Adaptivity

### v3

@@ Appendix A Regret Analysis of B-GLinCB
Additional Notation: We write c to denote absolute constant(s) that appears throughout our analysis. Our analysis also utilizes the following function
@@
We will show that ${\mathbf{H}}_{k}$ is an estimator of ${\mathbf{H}}^{*}_{k}$. Furthermore, we assume that the MLE estimator $\theta^{*}$ obtained by minimizing the log-loss objective always satisfies $\left\lVert\widehat{\theta}\right\rVert\leq S$. In case, that’s not true, one can use the non-convex projection described in Appendix E. The projected vector satisfies the same guarantees as described in the subsequent lemmas up to a multiplicative factor of $2$. Hence, the assumption $\left\lVert\widehat{\theta}\right\rVert\leq S$ is non-limiting.
@@
$=\sqrt{1+2RS}\left\lVert\left(\theta^{*}-\widehat{\theta}\right){\mathbf{G}}\right\rVert_{{\mathbf{G}}^{-1}}$
@@
$\leq\left\lVert y-y_{\varepsilon}\right\rVert_{2}\left\lVert\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-r_{j}\right)x_{j}\right\rVert_{{{\mathbf{H}}^{*}}^{-1}}+\langle y_{\varepsilon},{{\mathbf{H}}^{*}}^{-1/2}\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-r_{j}\right)x_{j}\rangle$
$\leq\varepsilon\left\lVert\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-r_{j}\right)x_{j}\right\rVert_{{{\mathbf{H}}^{*}}^{-1}}+\langle y_{\varepsilon},{{\mathbf{H}}^{*}}^{-1/2}\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-r_{j}\right)x_{j}\rangle$
@@
This result is derived directly from Lemma A.2 and the definition of (see ). By applying the lemma, we obtain
@@
Considering the definition of $\kappa$, we have $\dot{\mu}\left(\langle\mathbf{x},\theta^{*}\rangle\right)\geq\frac{1}{\kappa}$. This implies that ${\mathbf{H}}^{*}\succeq\frac{1}{\kappa}\mathbf{V}$ which in turn leads to the inequality $\left\lVert x\right\rVert_{{{\mathbf{H}}^{*}}^{-1}}\leq\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}$. ∎
@@
Via Lemma A.4 we have $|\langle x,\widehat{\theta}_{w}\rangle-\langle x,\theta^{*}\rangle|\leq\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\gamma(\lambda)$. Additionally, given that $\left\lVert\widehat{\theta}\right\rVert,\left\lVert\theta^{*}\right\rVert\leq S$ and $\left\lVert x\right\rVert\leq 1$ we also have $|\langle x,\widehat{\theta}_{w}\rangle-\langle x,\theta^{*}\rangle|\leq 2S$. Hence, we write
@@
Let $\widehat{\theta}_{k}$ be the estimator of $\theta^{*}$ calculated at the end of the $k^{th}$ batch. Then for any vector $x$ the following holds with probability greater than $1-\frac{\log{\log T}}{T^{2}}$ for every batch $k>=2$.
$|\langle x,\theta^{*}-\widehat{\theta}\rangle|\leq\left\lVert x\right\rVert_{{\mathbf{H}}_{k}^{-1}}\gamma{\left(\lambda\right)}$
@@
This result is a direct consequence of Lemma A.2 and the definition of $\gamma(\lambda)$ (see 7). According to the lemma, we have
$|\langle x,\theta^{*}-\widehat{\theta}_{w}\rangle|$ $\leq\left\lVert x\right\rVert_{{{\mathbf{H}}^{*}_{k}}^{-1}}\left\lVert{\theta^{*}-\widehat{\theta}_{w}}\right\rVert_{{\mathbf{H}}^{*}_{k}}$
@@
Let $x\in{\mathcal{X}}$ be the selected in any round of batch $k\geq 2$ in the algorithm, and let $x^{*}$ be the optimal arm in the arm set ${\mathcal{X}}$, i.e., $x^{*}=\operatorname*{arg\,max}_{x\in}{\mathcal{X}}$. With probability greater than $1-\frac{\log{\log{T}}}{T^{2}}$, the following inequality holds-
$\mu\left(\langle x^{*},\theta^{*}\rangle\right)-\mu\left(\langle x,\theta^{*}\rangle\right)\leq 6\phi(\lambda)\sum_{\begin{subarray}{c}y\in\{x,x^{*}\}\\ \widetilde{y}\in\{\widetilde{x},\widetilde{x}^{*}\}\end{subarray}}\left\lVert y\right\rVert_{\mathbf{V}^{-1}}\!\left\lVert\widetilde{y}\right\rVert_{{\mathbf{H}}_{k-1}}+2\gamma(\lambda)\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\left(\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}}+\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}}\right)$
@@
We begin by applying Taylor’s theorem, which yields the following for some $z\in\left[\langle x^{*},\theta^{*}\rangle,\langle x,\theta^{*}\rangle\right]$
@@
$\leq 2\dot{\mu}(z)\left(\left\lVert x^{*}\right\rVert_{{\mathbf{H}}_{k-1}}\gamma(\lambda)+\left\lVert x\right\rVert_{{\mathbf{H}}_{k-1}}\gamma(\lambda)\right)$ (via Lemma A.7)
$\leq 2\dot{\mu}(z)\gamma(\lambda)\left(\sqrt{\frac{\beta(x^{*})}{\dot{\mu}\left(\langle x^{*},\widehat{\theta}_{w}\rangle\right)}}\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}}+\sqrt{\frac{\beta(x)}{\dot{\mu}\left(\langle x,\widehat{\theta}_{w}\rangle\right)}}\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}}\right)$
$\leq 2\gamma(\lambda)\sqrt{\dot{\mu}(z)}\sqrt{\frac{\dot{\mu}(z)\beta(x^{*})}{\dot{\mu}\left(\langle x^{*},\widehat{\theta}_{w}\rangle\right)}}\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}}+2\sqrt{\dot{\mu}(z)}\gamma(\lambda)\sqrt{\frac{\dot{\mu}(z)\beta(x)}{\dot{\mu}\left(\langle x,\widehat{\theta}_{w}\rangle\right)}}\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}}$ (13)
@@
$\leq\exp\left(\min\left\{S,\frac{|\langle x,\theta^{*}\rangle-\langle x,\widehat{\theta}_{w}\rangle|+|\langle x,\theta^{*}\rangle-\langle x^{*},\theta^{*}\rangle|}{2}\right\}\right)\sqrt{\beta(x)}$ (since $z\in\left[\langle x^{*}_{t},\theta^{*}\rangle,\langle x,\theta^{*}\rangle\right]$ )
$\leq\exp\left(\min\left\{S,\frac{3\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\gamma\left(\lambda\right)+2\sqrt{\kappa}\left\lVert x^{*}\right\rVert_{\mathbf{V}^{-1}}\gamma\left(\lambda\right)}{2}\right\}\right)\sqrt{\beta(x)}$ (using Lemma A.4)
@@
$\sqrt{\dot{\mu}(z)}\left(\sqrt{\frac{\dot{\mu}(z)\beta(x^{*})}{\dot{\mu}\left(\langle x^{*},\widehat{\theta}_{w}\rangle\right)}}\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}}\right)$
$\leq 2\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\left(\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}}\exp\left(\min\left\{3S,3\sqrt{\kappa}\gamma\left(\lambda\right)\left(\left\lVert x\right\rVert_{\mathbf{V}^{-1}}+\left\lVert x^{*}\right\rVert_{\mathbf{V}^{-1}}\right)\right\}\right)\right)$
$\leq 6\frac{\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)\kappa}\ e^{3S}\ \gamma(\lambda)}{S}^{2}\left(\left\lVert x^{*}\right\rVert_{\mathbf{V}^{-1}}+\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\right)\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}}+2\gamma(\lambda)\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}}$
$\leq 6\frac{\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)\kappa}\ e^{3S}\ \gamma(\lambda)}{S}^{2}\left(\left\lVert x^{*}\right\rVert_{\mathbf{V}^{-1}}+\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\right)\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}}$
$\quad\quad+2\gamma(\lambda)\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}}$
$\leq 6\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\phi(\lambda)\left(\left\lVert x^{*}\right\rVert_{\mathbf{V}^{-1}}+\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\right)\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}}+2\gamma(\lambda)\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}}$
@@
$|\mu\left(\langle x^{*},\theta^{*}\rangle\right)-\mu\left(\langle x,\theta^{*}\rangle\right)|\leq$ $6\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\phi(\lambda)\left(\left\lVert x^{*}\right\rVert_{\mathbf{V}^{-1}}+\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\right)\left(\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}}+\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}}\right)$
$\quad\quad\quad+2\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\left(\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}}+\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}}\right)$
@@
$\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\sum_{\begin{subarray}{c}y\in\{x,x^{*}\}\\ \widetilde{y}\in\{\widetilde{x},\widetilde{x}^{*}\}\end{subarray}}\left\lVert y\right\rVert_{\mathbf{V}^{-1}}\left\lVert\widetilde{y}\right\rVert_{{\mathbf{H}}_{k-1}}$ $\leq 4\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\max_{x\in{\mathcal{X}}}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\max_{\mathbf{x}\in{\mathcal{X}}}\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}^{-}1}\right]$
@@
$\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\left(\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}}+\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}}\right)\right]$
$\leq 2\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\max_{x\in{\mathcal{X}}}\left\lVert x\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}\right]$
$\leq 2\min\left\{\sqrt{\kappa^{*}}\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\max_{x\in{\mathcal{X}}}\left\lVert x\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}\right],\sqrt{\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)\right]\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\left\lVert x\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}^{2}\right]}\right\}$ (using the definition of $\kappa^{*}$ for the first bound and Jensen for the second)
@@
A.2 Optimal Design Guarantees

### v4

@@ Appendix A Regret Analysis of B-GLinCB
A.1 Additional Notation
We write c to denote absolute constant(s) that appears throughout our analysis. Our analysis also utilizes the following function
@@ A.1 Additional Notation
We will show that ${\mathbf{H}}_{k}$ dominates ${\mathbf{H}}^{*}_{k}$ with high probability. Furthermore, we assume that the MLE estimator $\theta^{*}$ obtained by minimizing the log-loss objective always satisfies $\left\lVert\widehat{\theta}\right\rVert\leq S$. In case, that’s not true, one can use the non-convex projection described in Appendix E. The projected vector satisfies the same guarantees as described in the subsequent lemmas up to a multiplicative factor of $2$. Hence, the assumption $\left\lVert\widehat{\theta}\right\rVert\leq S$ is non-limiting.
A.2 Concentration Inequalities and Confidence Intervals
@@
$=\sqrt{1+2RS}\left\lVert{\mathbf{G}}\left(\theta^{*}-\widehat{\theta}\right)\right\rVert_{{\mathbf{G}}^{-1}}$
@@
$\leq\max_{y\in{\mathcal{B}}}\left\lVert y-y_{\varepsilon}\right\rVert_{2}\left\lVert\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-r_{j}\right)x_{j}\right\rVert_{{{\mathbf{H}}^{*}}^{-1}}+\langle y_{\varepsilon},{{\mathbf{H}}^{*}}^{-1/2}\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-r_{j}\right)x_{j}\rangle$
$\leq\max_{y\in{\mathcal{B}}}\varepsilon\left\lVert\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-r_{j}\right)x_{j}\right\rVert_{{{\mathbf{H}}^{*}}^{-1}}+\langle y_{\varepsilon},{{\mathbf{H}}^{*}}^{-1/2}\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-r_{j}\right)x_{j}\rangle$
@@
This result is derived directly from Lemma A.2 and the definition of $\gamma(\lambda)$ (see 9). By applying the lemma, we obtain
@@
Considering the definition of $\kappa$, we have $\dot{\mu}\left(\langle\mathbf{x},\theta^{*}\rangle\right)\geq\frac{1}{\kappa}$. This implies that ${\mathbf{H}}^{*}\succeq\frac{1}{\kappa}\mathbf{V}$  [^11] For logistic rewards, we note that instead of using the worst case bound of $\kappa$, one can make use of an upper bound on $S\coloneqq\left\lVert\theta\right\rVert$ as done in Mason et al. [2022] That is, we lower bound $\dot{\mu}\left(\langle x,\theta^{*}\rangle\right)\geq\dot{\mu}\left({S\left\lVert x\right\rVert}\right)$ and use ${\mathbf{H}}^{*}\succeq\sum_{t}\dot{\mu}\left({S\left\lVert x_{t}\right\rVert}\right)x_{t}x_{t}^{\intercal}$. which in turn leads to the inequality $\left\lVert x\right\rVert_{{{\mathbf{H}}^{*}}^{-1}}\leq\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}$. ∎
A.3 Preliminary Lemmas
@@
Via Corollary A.4 we have $|\langle x,\widehat{\theta}_{w}\rangle-\langle x,\theta^{*}\rangle|\leq\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\gamma(\lambda)$. Additionally, given that $\left\lVert\widehat{\theta}\right\rVert,\left\lVert\theta^{*}\right\rVert\leq S$ and $\left\lVert x\right\rVert\leq 1$ we also have $|\langle x,\widehat{\theta}_{w}\rangle-\langle x,\theta^{*}\rangle|\leq 2S$. Hence, we write
@@
Let $\widehat{\theta}_{k}$ be the estimator of $\theta^{*}$ calculated at the end of the $k^{th}$ batch. Then for any vector $x$ the following holds with probability greater than $1-\frac{\log{\log T}}{T^{2}}$ for every batch $k\geq 2$.
$|\langle x,\theta^{*}-\widehat{\theta}_{k}\rangle|\leq\left\lVert x\right\rVert_{{\mathbf{H}}_{k}^{-1}}\gamma{\left(\lambda\right)}$
@@
This result is a direct consequence of Lemma A.2 and the definition of $\gamma(\lambda)$ (see 9). According to the lemma, we have
$|\langle x,\theta^{*}-\widehat{\theta}_{k}\rangle|$ $\leq\left\lVert x\right\rVert_{{{\mathbf{H}}^{*}_{k}}^{-1}}\left\lVert{\theta^{*}-\widehat{\theta}_{k}}\right\rVert_{{\mathbf{H}}^{*}_{k}}$
@@
Let $x\in{\mathcal{X}}$ be the selected in any round of batch $k\geq 2$ in the algorithm, and let $x^{*}$ be the optimal arm in the arm set ${\mathcal{X}}$, i.e., $x^{*}=\operatorname*{arg\,max}_{x\in{\mathcal{X}}}\mu\left(\langle x,\theta^{*}\rangle\right)$. With probability greater than $1-\frac{\log{\log{T}}}{T^{2}}$, the following inequality holds-
$\mu\left(\langle x^{*},\theta^{*}\rangle\right)-\mu\left(\langle x,\theta^{*}\rangle\right)\leq 6\phi(\lambda)\sum_{\begin{subarray}{c}y\in\{x,x^{*}\}\\ \widetilde{y}\in\{\widetilde{x},\widetilde{x}^{*}\}\end{subarray}}\left\lVert y\right\rVert_{\mathbf{V}^{-1}}\!\left\lVert\widetilde{y}\right\rVert_{{\mathbf{H}}_{k-1}}+2\gamma(\lambda)\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\left(\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}+\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}\right)$
@@
We begin by applying Taylor’s theorem, which yields the following for some $z$ between $\langle x,\theta^{*}\rangle$ and $\langle x^{*},\theta^{*}\rangle$
@@
$\leq 2\dot{\mu}(z)\left(\left\lVert x^{*}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}\gamma(\lambda)+\left\lVert x\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}\gamma(\lambda)\right)$ (via Corollary A.7)
$\leq 2\dot{\mu}(z)\gamma(\lambda)\left(\sqrt{\frac{\beta(x^{*})}{\dot{\mu}\left(\langle x^{*},\widehat{\theta}_{w}\rangle\right)}}\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}+\sqrt{\frac{\beta(x)}{\dot{\mu}\left(\langle x,\widehat{\theta}_{w}\rangle\right)}}\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}\right)$
$\leq 2\gamma(\lambda)\sqrt{\dot{\mu}(z)}\sqrt{\frac{\dot{\mu}(z)\beta(x^{*})}{\dot{\mu}\left(\langle x^{*},\widehat{\theta}_{w}\rangle\right)}}\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}+2\sqrt{\dot{\mu}(z)}\gamma(\lambda)\sqrt{\frac{\dot{\mu}(z)\beta(x)}{\dot{\mu}\left(\langle x,\widehat{\theta}_{w}\rangle\right)}}\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}$ (15)
@@
$\leq\exp\left(\min\left\{S,\frac{|\langle x,\theta^{*}\rangle-\langle x,\widehat{\theta}_{w}\rangle|+|\langle x,\theta^{*}\rangle-\langle x^{*},\theta^{*}\rangle|}{2}\right\}\right)\sqrt{\beta(x)}$ (since $z\in\left[\langle x,\theta^{*}\rangle,\langle x^{*}_{t},\theta^{*}\rangle\right]$ )
$\leq\exp\left(\min\left\{S,\frac{3\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\gamma\left(\lambda\right)+2\sqrt{\kappa}\left\lVert x^{*}\right\rVert_{\mathbf{V}^{-1}}\gamma\left(\lambda\right)}{2}\right\}\right)\sqrt{\beta(x)}$ (using Lemma A.4 and the elimination criteria)
@@
$\sqrt{\dot{\mu}(z)}\left(\sqrt{\frac{\dot{\mu}(z)\beta(x^{*})}{\dot{\mu}\left(\langle x^{*},\widehat{\theta}_{w}\rangle\right)}}\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}\right)$
$\leq 2\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\left(\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}\exp\left(\min\left\{3S,3\sqrt{\kappa}\gamma\left(\lambda\right)\left(\left\lVert x\right\rVert_{\mathbf{V}^{-1}}+\left\lVert x^{*}\right\rVert_{\mathbf{V}^{-1}}\right)\right\}\right)\right)$
$\leq 6\frac{\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)\kappa}\ e^{3S}\ \gamma(\lambda)}{S}^{2}\left(\left\lVert x^{*}\right\rVert_{\mathbf{V}^{-1}}+\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\right)\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}+2\gamma(\lambda)\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}$ (via Claim A.8)
$\leq 6\frac{\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)\kappa}\ e^{3S}\ \gamma(\lambda)}{S}^{2}\left(\left\lVert x^{*}\right\rVert_{\mathbf{V}^{-1}}+\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\right)\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}$
$\quad\quad+2\gamma(\lambda)\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}$
$\leq 6\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\phi(\lambda)\left(\left\lVert x^{*}\right\rVert_{\mathbf{V}^{-1}}+\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\right)\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}+2\gamma(\lambda)\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}$
@@
$|\mu\left(\langle x^{*},\theta^{*}\rangle\right)-\mu\left(\langle x,\theta^{*}\rangle\right)|\leq$ $6\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\phi(\lambda)\left(\left\lVert x^{*}\right\rVert_{\mathbf{V}^{-1}}+\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\right)\left(\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}+\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}\right)$
$\quad\quad\quad+2\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\left(\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}+\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}\right)$
@@
$\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\sum_{\begin{subarray}{c}y\in\{x,x^{*}\}\\ \widetilde{y}\in\{\widetilde{x},\widetilde{x}^{*}\}\end{subarray}}\left\lVert y\right\rVert_{\mathbf{V}^{-1}}\left\lVert\widetilde{y}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}$ $\leq 4\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\max_{x\in{\mathcal{X}}}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\max_{\mathbf{x}\in{\mathcal{X}}}\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}\right]$
@@
$\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\left(\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}+\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}\right)\right]$
$\leq 2\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\max_{x\in{\mathcal{X}}}\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}\right]$
$\leq 2\min\left\{\sqrt{\frac{1}{\kappa^{*}}}\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\max_{x\in{\mathcal{X}}}\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}\right],\sqrt{\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)\right]\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}^{2}\right]}\right\}$ (using the definition of $\kappa^{*}$ for the first bound and Jensen for the second)
@@
A.5 Optimal Design Guarantees